OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1). parallel: y = -3x + 6 perpendicular: y = one over three x + two and two thirds parallel: y = -3x + 10 perpendicular: y = one over three x parallel: y = negative one over three x perpendicular: y = -3x parallel: y = one over three x + 1 perpendicular: y = 3x - 1

OpenStudy (jdoe0001):

keep in mind that parallel lines will have the same slope as y = -3x-1 and perpendicular lines will have a NEGATIVE RECIPROCAL slopes of that notice the slope of y = -3x-1 \(\large y={\color{brown}{ -3}}x-1 \begin{cases} parallel\\ \hline\\ slope\to {\color{brown}{ -3}}\\ perpendicular\\ \hline\\ {\color{brown}{ -3}}\quad negative\to {\color{brown}{ +3}}\quad reciprocal\to \frac{1}{{\color{brown}{ +3}}} \end{cases}\)